evaluate the following integral
∫ sin 2x / a^2 + b^2 sin^2 (x) dx

Grade:12

1 Answers

Sanjeev Kumar
9 years ago
let a^2 + b^2 sin^2 (x) = t
b^2 sin 2x dx = dt
sin 2x dx = dt/b2
equating the the above value in 
∫sin 2x / a^2 + b^2 sin^2 (x) dx
we get 
∫dt/b2t
= 1/b∫dt/t
= 1/b2 log t
= 1/b2 log (a^2 + b^2 sin^2 (x))